Orange, Inc. has identified the following cost drivers for its expected overhead costs for the year:

Overhead Item Expected Cost Cost Driver Expected Quantity
Setup costs $ 50,000 Number of setups 250
Ordering costs 30,000 Number of orders 1,500
Maintenance 100,000 Machine hours 2,000
Power 20,000 Kilowatt hours 4,000
Total Overhead $ 200,000

Total direct labor hours budgeted = 2,000 hours.

The following actual data applies to one of the products completed during the year:

Product X
Direct materials $ 5,000 Number of setups 5
Direct labor $ 3,000 Number of orders 50
Units completed 100 Machine hours 50
Direct labor hours 100 Kilowatt hours 500

If Orange, Inc. uses direct labor hours to assign overhead, the unit product cost for Product X will be:

Multiple Choice

  • $70.00.

  • $60.00.

  • $180.00.

  • $90.00.

  • $80.00
Answer #1

Overhead rate = Estimated overhead/Budgeted direct labor hours

= 200,000/2,000

= 100 per dlh

Unit product cost = (5,000/100) + (3,000/100) + (100*100)/100

= 50 + 30 + 100

= 180